OpenStudy (anonymous):

Compare: three-fourths to nine-twelfths Choose one answer. a. less than b. greater than c. equal to d. None of the above

OpenStudy (anonymous):

You again :D

OpenStudy (anonymous):

this would be the same as the last one, convert 3/4 to x/12

OpenStudy (anonymous):

3*4=12 and 3*3=9 so 3/4=9/12

OpenStudy (anonymous):

so they would be equal

OpenStudy (anonymous):

So c?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you so much :D medal and a fan :D